### v2

# Efficient Anomaly Detection with Budget Annotation Using Semi-Supervised Residual Transformer

Abstract
Anomaly Detection (AD) is challenging as usually only the normal samples are seen during training and the detector needs to discover anomalies on-the-fly. The recently proposed deep learning based approaches are able to alleviate this problem but there is still a long way to go in developing anomaly detectors that meet most real-world industrial applications. On the other hand, in some particular AD tasks, a few anomalous samples are labeled manually for achieving higher accuracy. However, this performance gain is at the cost of considerable annotation efforts, which can be intractable in many practical scenarios.
In this work, the above two problems are addressed in a unified framework. Firstly, inspired by the success of the patch-matching-based AD algorithms, we train a sliding vision transformer model over the residuals generated by a novel position-constrained patch-matching. Secondly, the conventional pixel-wise segmentation problem is cast into a block-wise classification problem. Thus the sliding transformer can attain even higher accuracy with much less annotation labor. Thirdly, to further reduce the labeling cost, we propose to label the anomalous regions using only bounding boxes. The unlabeled regions caused by the weak labels are effectively exploited using a highly-customized semi-supervised learning scheme equipped with two novel data augmentation methods. The proposed method, termed “Semi-supervised RESidual Transformer” or “SemiREST” in short, outperforms all recent state-of-the-art approaches using all the evaluation metrics in both the unsupervised and supervised scenarios. On the popular MVTec-AD dataset, our SemiREST algorithm obtains the Average Precision (AP) of (vs. previous best result of AP ) in the unsupervised condition and AP (vs. previous best result of AP ) for supervised anomaly detection. Notably, with the bounding-box-based semi-supervisions, SemiREST still outperforms recent representative methods with full supervision ( AP vs. previous best AP ) on MVTec-AD. Similar precision advantages are also observed on the other two well-known AD datasets, i.e., BTAD, and KSDD2. Overall, the proposed SemiREST generates new records of AD performances while at a remarkably low annotation cost. It is not only cheaper in annotation but also performs better than most recent methods. The code of this work is available at: https://github.com/BeJane/Semi_REST
Keywords:
Anomaly detection Vision Transformer Semi-supervised learning.1 Introduction
Product quality control is crucial in many manufacturing processes. Manual inspection is expensive and unreliable considering the limited time budget for inspection on a running assembly line. As a result, automatic defect inspection is in great demand in modern manufacturing industries (Niu et al., 2021; Ni et al., 2021; Tao et al., 2022; Cao et al., 2023b). Given training samples with sufficient supervision, it is straightforward to perform defect detection using off-the-shelf detection or segmentation algorithms (Cheng et al., 2018; Wang et al., 2022a, c, b). Unfortunately, in most practical cases, one can only obtain much fewer “anomaly” samples than the normal ones and the anomaly pattern varies dramatically over different samples. As a result, the defect detection task is often cast into an Anomaly Detection (AD) problem (Bergmann et al., 2019; Mishra et al., 2021), with only normal instances available for training.
As a result, the majority of the literature on anomaly detection focuses on learning the “one-class” patterns from the anomaly-free training set. The most straightforward way of realizing AD is to treat the normal image patches as one class while the anomalous ones as outliers (Ruff et al., 2018; Yi and Yoon, 2020; Liznerski et al., 2021; Massoli et al., 2021; Defard et al., 2021; Zhang et al., 2022; Chen et al., 2022). Some researchers propose to localize the anomalies by comparing the test image patches with the normal references which are directly captured from the training set (Roth et al., 2022; Kim et al., 2022; Bae et al., 2022; Saiku et al., 2022; Zhu et al., 2022; Xie et al., 2023) or reconstructed based on the normal samples (Zong et al., 2018; Dehaene and Eline, 2020; Shi et al., 2021; Hou et al., 2021; Wu et al., 2021). Besides, the distillation-based methods (Salehi et al., 2021; Deng and Li, 2022; Bergmann et al., 2022; Zhang et al., 2023b) and latent image registration (Huang et al., 2022; Liu et al., 2023a) are also proposed to enhance the outlier effect of anomalies.
Despite the mainstream “unsupervised” fashion, in (Ding et al., 2022; Zhang et al., 2023a; Yao et al., 2023), anomaly detectors are trained in a “few-shot” manner and unsurprisingly higher performances are achieved, compared with their “zero-shot” counterparts. In addition, the traditional surface defect detection tasks (Tabernik et al., 2020; Huang et al., 2020; Zhang et al., 2021b; Božič et al., 2021) are mostly solved in a fully-supervised version. Considering that in practice, the detection algorithm needs to “warm-up” for a certain duration along with the running assembly line, the abnormal samples are not difficult to obtain in this scenario. And thus this “supervised” setting is actually valid.
In this paper, we claim that compared with the lack of training anomalies, a more realistic problem to address is the limited time budget for labeling the newly-obtained anomalous images. Accordingly, we propose a novel AD algorithm that simultaneously enjoys high detection robustness and low annotation cost. The high-level concept of the proposed algorithm is illustrated in Figure 1. Compared with the conventional pixel-wise supervision, the proposed method can utilize weak labels, i.e., the block labels or bounding-box labels (as shown in the bottom part of Figure 1) to get even higher AD performances.
In specific, we consider the AD task as a block-wise classification problem (as shown in the right-middle part of Figure 1) and thus one only needs to label hundreds of anomaly blocks instead of thousands of anomaly pixels on a defective image. The patch-matching residual is generated for each block constrained by the block’s position-code. Then the blocks are classified by a sliding Swin Transformer (Liu et al., 2021) that receives the patch-matching residuals as its tokens. The final anomaly score of a block is obtained via a bagging process over all the Swin transformers whose input involves this block. Thanks to the novel features and the bagging strategy, the proposed algorithm performs better than those with pixel-level supervision. More aggressively, we propose to further reduce the annotation cost by labeling the anomaly regions with only bounding boxes. As can be seen in the right-bottom part of Figure 1, the boxes cover all the anomaly regions of the image so the outside blocks (shown in green) can be directly used as negative samples. On the other hand, the identities of the inside blocks (shown in yellow) are not given and seem useless in the conventional learning paradigm. In this work, nonetheless, the unlabeled blocks are utilized effectively via a modified Mix-Match (Berthelot et al., 2019b) algorithm that is smartly customized for the transformer-based anomaly detector. The productive utilization of the unlabeled information reduces the performance drop from full supervision.
We term the proposed AD algorithm “Semi-supervised RESidual Transformer” or “SemiREST” in short. Our experiments in this paper verify the superiority of the proposed method: SemiREST outperforms all the state-of-the-art AD algorithms on the three well-known datasets (MVTec-AD (Bergmann et al., 2019), BTAD (Mishra et al., 2021) and KSDD2 (Božič et al., 2021)) in both the mainstream “unsupervised” setting and the “few-shot” setting. In addition, with only bounding-box annotations, the SemiREST also beats all the SOTA methods with fully-supervised pixel labels with all the employed evaluation metrics. In summary, our main contribution is three fold, as listed below.
Cheaper in terms of annotation: From a realistic perspective, we suggest that, compared with the “one-class” compatibility, a more desirable property for AD algorithms is the high efficiency of using annotation information. Accordingly, this work proposes two types of low-cost annotations: the block-wise labels and the bounding box labels. To the best of our knowledge, it is the first time that the block-wise label is used for anomaly detection and the usage of the bounding-box label in the context of semi-supervised learning is also creative in the literature of AD.
Better: We design the SemiREST algorithm based on a modified Swin-Transformer (Liu et al., 2021) and the patch-matching residuals generated with position-code constraints. The proposed algorithm consistently surpasses the SOTA methods by large margins, on the three most acknowledged datasets and in all the supervision conditions.
Cheaper and Better: Given the cheap bounding-box labels, SemiREST can effectively utilize the unlabeled features thanks to the proposed ResMixMatch algorithm, which is highly customized from MixMatch (Berthelot et al., 2019b) for the scenario of residual-based anomaly detection. Simply using these lightweight annotations, SemiREST still outperforms all the comparing SOTA methods trained with fully annotated pixel-wise labels.
2 Related Work
2.1 Recently Proposed Anomaly Detection Algorithms
During the past few years, a great amount of effort has been invested in developing better AD algorithms. The yielded AD algorithms could be mainly divided into three categories.
1.
One-class classification
Following the pioneering works (Scholkopf et al., 2000; Schölkopf et al., 2001), a large part of AD literature focus on modeling normal images or patches as a single class while the abnormal ones are detected as outliers (Ruff et al., 2018; Yi and Yoon, 2020; Liznerski et al., 2021; Massoli et al., 2021; Defard et al., 2021; Zhang et al., 2022; Chen et al., 2022). Normalizing Flow (Dinh et al., 2014, 2016; Kingma and Dhariwal, 2018) based AD algorithms propose to further map the normal class to fit a Gaussian distribution in the feature space (Rudolph et al., 2021; Yu et al., 2021; Tailanian et al., 2022; Lei et al., 2023). -
2.
Patch matching
By observing that most normal samples can be effectively reconstructed by other normal ones, a number of researchers have tried to conduct the reconstruction in various spaces for the test images or image patches and estimate the anomaly scores based on the reconstruction residuals (Zong et al., 2018; Dehaene and Eline, 2020; Shi et al., 2021; Hou et al., 2021; Wu et al., 2021). In addition to the reconstruction, some more sophisticated anomaly detectors (Li et al., 2021; Zavrtanik et al., 2021; Yang et al., 2023) also learn discriminative models with “pseudo anomalies” for higher accuracy. At the other end of the spectrum, however, the seminar work PatchCore (Roth et al., 2022) shows that a well-designed nearest-neighbor matching process can already achieve sufficiently good detection accuracies. -
3.
Distillation
To obtain distinct responses to anomaly inputs, distillation-based approaches (Salehi et al., 2021; Deng and Li, 2022; Bergmann et al., 2022; Zhang et al., 2023b) distill a “teacher network” (which is pre-trained on a large but “neutral” dataset) into a “student network” only based on normal samples. Then the anomaly score can be calculated according to the response differences between the two networks.
2.2 PathCore Algorithm and Its Variants
As a simple and typical example of the patch-matching-based AD methods, PatchCore algorithm (Roth et al., 2022) proposes the coreset-subsampling algorithm to build a “memory bank” of patch features, which are obtained via smoothing the neutral deep features pre-learned on ImageNet (Deng et al., 2009; Russakovsky et al., 2015). The anomaly score is then calculated based on the Euclidean distance between the test patch feature and its nearest neighbor in the “memory bank”. Despite the simplicity, PatchCore performs dramatically well on the MVTec-AD dataset (Bergmann et al., 2019).
The good performance of PatchCore encourages researchers to develop better variants based on it. The PAFM algorithm (Kim et al., 2022) applies patch-wise adaptive coreset sampling to improve the speed. (Bae et al., 2022) introduces the position and neighborhood information to refine the patch-feature comparison. Graphcore (Xie et al., 2023) utilizes graph representation to customize PatchCore for the few-shot setting. (Saiku et al., 2022) modifies PatchCore by compressing the memory bank via k-means clustering. (Zhu et al., 2022) combines PatchCore (Roth et al., 2022) and Defect GAN (Zhang et al., 2021a) for higher AD performances. Those variants, though achieving slightly better performances, all fail to notice the potential value of the intermediate information generated by the patch-matching. In this work, we use the matching residuals as the input tokens of our transformer model. The individual and the mutual information of the residuals are effectively exploited and new SOTA performances are then obtained.
2.3 Swin Transformer for Anomaly Detection
As a variation of Vision Transformer (ViT) (Dosovitskiy et al., 2021), Swin Transformer (Liu et al., 2021, 2022) proposed a hierarchical Transformer with a shifted windowing scheme, which not only introduces several visual priors into Transformer but also reduces computation costs. Swin Transformer and its variants illustrate remarkable performances in various computer vision tasks, such as semantic segmentation (Huang et al., 2021; Hatamizadeh et al., 2022; Cao et al., 2023a), instance segmentation (Dong et al., 2021; Li et al., 2022) and object detection (Xu et al., 2021; Dai et al., 2021; Liang et al., 2022).
In the literature on anomaly detection, some recently proposed methods also employ Swin Transformer as the backbone network. (Üzen et al., 2022) develops a hybrid structure decoder that combines convolution layers and the Swin Transformer. (Gao et al., 2022) improves the original shifted windowing scheme of the Swin Transformer for surface defect detection. Despite the success of Swin Transformer models in other domains, the Swin-Transformer-based AD algorithms could hardly outperform the SOTA methods on most acknowledged datasets such as (Bergmann et al., 2019), (Mishra et al., 2021) and (Božič et al., 2021).
In this paper, we successfully tame the Swin Transformer for the small training sets of AD problems by introducing a series of novel modifications.
2.4 MixMatch and Weak Labels Based on Bounding Boxes
Semi-supervised Learning (SSL) is always attractive since it can save massive labeling labor. Many efforts have been devoted to utilizing the information from the unlabeled data (Zhu and Goldberg, 2009; Berthelot et al., 2019b, a; Sohn et al., 2020; Wang et al., 2023), mainly focusing on the generation of high-quality pseudo labels. Inspired by the seminar work (Zhang et al., 2018; Yun et al., 2019) for data augmentation, MixMatch proposes a single-loss SSL method that relies on a smart fusion process between labeled and unlabeled samples. Thus, MixMatch enjoys high accuracy and a relatively simple training scheme.
On the other hand, in the literature on semantic segmentation, bounding boxes are usually used as weak supervision to save labeling costs. (Hsu et al., 2019) exploits the tightness prior to the bounding boxes to generate the positive and negative bags for multiple instance learning (MIL), which is integrated into a fully supervised instance segmentation network. (Kervadec et al., 2020) integrates the tightness prior and a global background emptiness constraint derived from bounding box annotations into a weak semantic segmentation of medical images. (Lee et al., 2021) propose a bounding box attribution map (BBAM) to produce pseudo ground-truth for weakly supervised semantic and instance segmentation.
In this work, within the block-wise classification framework, MixMatch is smartly tailored to exploit the information of unlabeled blocks which are brought by the weak supervision of bounding boxes. This combination of the novel semi-supervised learning scheme and the bounding box labels is remarkably effective according to the experiment results and also novel in the literature, to our best knowledge.
3 Experiments
In this section, extensive experiments are conducted to evaluate the proposed method, compared with a comprehensive collection of SOTA methods (Yang et al., 2023; Zavrtanik et al., 2021; Deng and Li, 2022), (Ristea et al., 2022; Liu et al., 2023a, b; Zhang et al., 2023b; Lei et al., 2023; Gudovskiy et al., 2022; Tien et al., 2023; Zhang et al., 2023a; Yao et al., 2023; Pang et al., 2021; Ding et al., 2022), on three well-acknowledged benchmarks, namely, the MVTec-AD (Bergmann et al., 2019) dataset, the BTAD (Mishra et al., 2021) dataset and the KolektorSDD2 dataset (Božič et al., 2021), respectively.
3.1 Three levels of supervision
First of all, let us clarify the experiment settings for the three types of supervision as we introduced in LABEL:subsec:supervision.
In the unsupervised setting (Un), only normal data can be accessed during training. The involved algorithms are evaluated on the entire test set, including both normal and anomalous images.
In the supervised setting (Sup), we randomly draw anomalous images with block-wise annotations from various types of defects to the train set and remove them from the test set. We also involve the original anomaly-free training samples to learn the model in the setting. Note that this setting strictly follows the data splitting principle in (Zhang et al., 2023a) and (Ding et al., 2022). The proposed methods only used the block-wise labels which are obtained by using Equation LABEL:equ:block_label for this supervision condition.
The semi-supervised setting (Semi) is the proposed supervision condition in this work. The same anomalous images are used as training samples while the block labels of them are determined by the bounding-boxes which are defined in Equation LABEL:equ:block_label_semi. Note that in this case, all the blocks are annotated as either unknown or normal.
3.2 Implementation details
The layer- and layer- feature maps of the WideResnet- model (Zagoruyko and Komodakis, 2016) (pre-trained on ImageNet-1K) are concatenated with a proper scale adaptation and the yielded “hyper-columns” are smoothed via average pooling in the neighborhoods. We set to generate PCFs. For MVTec-AD (Bergmann et al., 2019) and BTAD (Mishra et al., 2021) datasets, we subsample of the PCFs to get the memory banks , while the subsampling ratio for KolektorSDD2 (Božič et al., 2021) is considering the much larger training set of normal images. Our Swin Transformer model consists of blocks with patch csize , the window size is and the number of heads is set to . We train the unsupervised and supervised models from scratch and employ the unsupervised model to initialize the weights of the semi-supervised model as Section LABEL:subsec:mixmatch introduces. The parameters of the focal loss , , and are set to , , and respectively. The sliding window slides over the PCR tensor with a step and the window size is . and are set to and respectively. To speed up training, we sample of sliding windows in normal images, simulated defective images, and true anomalous images at each training iteration. To compare the final prediction map with the ground-truth label map, it is firstly upscaled to the same size as the ground-truth via bilinear interpolation and then smoothed using a Gaussian of kernel with , as it is done in (Roth et al., 2022).
The values of , , , , , and vary among different supervision conditions:
As to the semi-supervised learning, the hyperparameters , , and are set to , , and respectively. The random noise for agumentation is set to . Following the MixcMatch algorithm (Berthelot et al., 2019b), we linearly ramp up the unlabeled loss weight to ( for BTAD) over the first steps of training.
3.3 Evaluation methods
In this work, the involved AD algorithms are measured comprehensively by three popular threshold-indeendent metrics: Pixel-AUROC, PRO (Bergmann et al., 2020) (per region overlap) and AP (Zavrtanik et al., 2021) (average precision). In specific, Pixel-AUROC is the area under the receiver operating characteristic curve at the pixel level. It is the most popular AD measuring method while failing to reflect the real performance difference between algorithms when a serious class imbalance exists. The PRO score, on the contrary, focuses on the anomaly pixels and treats the AD performance on each anomaly region equally. Consequently, the PRO metric is more robust to the class imbalance which is actually a common situation in most AD benchmarks. The AP (average precision) metric (Zavrtanik et al., 2021), as a conventional metric for semantic segmentation, is frequently adopted in recently proposed AD algorithms (Zavrtanik et al., 2021; Zhang et al., 2023a). It reflects the anomaly detection performance from a pixel-level perspective.
3.4 Results on MVTec-AD
MVTec-AD (Bergmann et al., 2019) is the most popular AD dataset with high-resolution color images belonging to texture categories and object categories. Each category contains a train set with only normal images and a test set with various kinds of defects as well as defect-free images. We conduct the experiments on this dataset within all three supervision conditions.
The unsupervised AD results of the comparing algorithms on MVTec-AD (Bergmann et al., 2019) are shown in Table 1. As shown in the table, our method achieves the highest average AP, average PRO and average pixel AUROC, for both texture and object categories and outperforms the unsupervised SOTA by , and respectively. In specific, SemiREST ranks first on ( out of ) categories with AP metric and the “first-ranking” ratios for the PRO and Pixel-AUROC are and .
In addition, Table 2 illustrates that with full supervision, SemiREST still ranks first for the average AD performance evaluated by using all three metrics. In particular, our method outperforms the supervised SOTAs by on AP, on PRO and on Pixel-AUROC. The “first-ranking” ratios of SemiREST in the supervised scenario are , and one AP, PRO and Pixel-AUROC respectively.
Table 2 also reports the performances of our method with the bounding-box-determined semi-supervisions. It can be seen that the semi-supervised SemiREST performs very similarly to the fully-supervised SemiREST, thanks to the effective usage of the unlabeled blocks. In addition, even with much less annotation information, the semi-supervised SemiREST still beats the fully-supervised SOTA methods by large margins ( for AP, for PRO and for Pixel-AUROC).
It is interesting to see that with only synthetic defective samples, the unsupervised SemiREST remains superior to the supervised SOTA, with the AP and Pixel-AUROC metrics (see Table 1 and see Table 2). The proposed algorithm illustrates remarkably high generalization capacities.
Readers can also find the qualitative results of the proposed method compared with other SOTA algorithms in Figure 4.
| Category | MemSeg | PatchCore | DRAEM | RD | SSPCAB | NFAD | DMAD | SimpleNet | DeSTSeg | PyramidFlow | RD++ | Ours (Un) |
| (Yang et al., 2023) | (Roth et al., 2022) | (Zavrtanik et al., 2021) | (Deng and Li, 2022) | (Ristea et al., 2022) | (Yao et al., 2023) | (Liu et al., 2023a) | (Liu et al., 2023b) | (Zhang et al., 2023b) | (Lei et al., 2023) | (Tien et al., 2023) | ||
| Carpet | 64.3/91.2/99.2 | 64.1/95.1/99.1 | 53.5/92.9/95.5 | 56.5/95.4/98.9 | 48.6/86.4/92.6 | 74.1/98.2/99.4 | 63.8/95.9/99.0 | 44.1/92.0/97.7 | 72.8//96.1 | /97.2/97.4 | /97.7/99.2 | 84.2/98.7/99.6 |
| Grid | 41.6/96.2/99.1 | 30.9/93.6/98.8 | 65.7/98.3/99.7 | 15.8/94.2/98.3 | 57.9/98.0/99.5 | 51.9/97.9/99.3 | 47.0/97.3/99.2 | 39.6/94.6/98.7 | 61.5//99.1 | /94.3/95.7 | /97.7/99.3 | 65.5/97.9/99.5 |
| Leather | 73.2/99.3/99.8 | 45.9/97.2/99.3 | 75.3/97.4/98.6 | 47.6/98.2/99.4 | 60.7/94.0/96.3 | 70.1/99.4/99.7 | 53.1/98.0/99.4 | 48.0/97.5/99.2 | 75.6//99.7 | /99.2/98.7 | /99.2/99.4 | 79.3/99.4/99.8 |
| Tile | 95.6/98.5/99.6 | 54.9/80.2/95.7 | 92.3/98.2/99.2 | 54.1/85.6/95.7 | 96.1/98.1/99.4 | 63.0/91.8/96.7 | 56.5/84.3/95.8 | 63.5/78.3/93.9 | 90.0//98.0 | /97.2/97.1 | /92.4/96.6 | 96.4/98.5/99.7 |
| Wood | 81.8/94.5/98.0 | 50.0/88.3/95.0 | 77.7/90.3/96.4 | 48.3/91.4/95.8 | 78.9/92.8/96.5 | 62.9/95.6/96.9 | 45.5/89.3/94.8 | 48.8/83.9/93.9 | 81.9//97.7 | /97.9/97.0 | /93.3/95.8 | 79.4/96.5/97.7 |
| Average | 71.3/95.9/99.1 | 49.2/90.9/97.6 | 72.9/95.4/97.9 | 44.5/93.0/97.6 | 68.4/93.9/96.9 | 64.4/96.6/98.4 | 53.2/93.0/97.6 | 48.8/89.3/96.7 | 76.4//98.1 | /97.2/97.2 | /96.1/98.1 | 81.0/98.2/99.3 |
| Bottle | 89.5/96.9/99.3 | 77.7/94.7/98.5 | 86.5/96.8/99.1 | 78.0/96.3/98.8 | 89.4/96.3/99.2 | 77.9/96.6/98.9 | 79.6/96.4/98.8 | 73.0/91.5/98.0 | 90.3//99.2 | /95.5/97.8 | /97.0/98.8 | 94.1/98.6/99.6 |
| Cable | 74.3/90.3/98.0 | 66.3/93.2/98.4 | 52.4/81.0/94.7 | 52.6/94.1/97.2 | 52.0/80.4/95.1 | 65.7/95.9/98.0 | 58.9/92.2/97.9 | 69.3/89.7/97.5 | 60.4//97.3 | /90.3/91.8 | /93.9/98.4 | 81.1/95.3/99.1 |
| Capsule | 56.8/95.7/99.2 | 44.7/94.8/99.0 | 49.4/82.7/94.3 | 47.2/95.5/98.7 | 46.4/92.5/90.2 | 58.7/96.0/99.2 | 42.2/91.6/98.1 | 44.7/92.8/98.9 | 56.3//99.1 | /98.3/98.6 | /96.4/98.8 | 57.2/96.9/98.8 |
| Hazelnut | 76.8/91.8/99.2 | 53.5/95.2/98.7 | 92.9/98.5/99.7 | 93.4/98.2/99.7 | 60.0/95.2/98.6 | 65.3/97.6/98.6 | 63.4/95.9/99.1 | 48.3/92.2/97.6 | 88.4//99.6 | /98.1/98.1 | /96.3/99.2 | 87.8/96.1/99.6 |
| Metal nut | 95.2/96.8/99.2 | 86.9/94.0/98.3 | 96.3/97.0/99.5 | 78.6/94.9/97.3 | 94.7/97.7/99.4 | 76.6/94.9/97.7 | 79.0/94.2/97.1 | 92.6/91.3/98.7 | 93.5//98.6 | /91.4/97.2 | /93.0/98.1 | 96.6/97.5/99.5 |
| Pill | 89.1/97.9/99.5 | 77.9/95.0/97.8 | 48.5/88.4/97.6 | 76.5/96.7/98.1 | 48.3/89.6/97.2 | 72.6/98.1/98.0 | 79.7/96.9/98.5 | 80.1/93.9/98.5 | 83.1//98.7 | /96.1/96.1 | /97.0/98.3 | 85.9/98.4/99.2 |
| Screw | 52.2/95.6/99.1 | 36.1/97.1/99.5 | 58.2/95.0/97.6 | 52.1/98.5/99.7 | 61.7/95.2/99.0 | 47.4/96.3/99.2 | 47.9/96.5/99.3 | 38.8/95.2/99.2 | 58.7//98.5 | /94.7/94.6 | /98.6/99.7 | 65.9/97.9/99.7 |
| Toothbrush | 60.9/96.5/99.4 | 38.3/89.4/98.6 | 44.7/85.6/98.1 | 51.1/92.3/99.1 | 39.3/85.5/97.3 | 38.8/92.3/98.7 | 71.4/91.5/99.3 | 51.7/88.7/98.6 | 75.2//99.3 | /97.9/98.5 | /94.2/99.1 | 74.5/96.2/99.5 |
| Transistor | 82.1/92.8/97.2 | 66.4/92.4/96.3 | 50.7/70.4/90.9 | 54.1/83.3/92.3 | 38.1/62.5/84.8 | 56.0/82.0/94.0 | 58.5/85.2/94.1 | 69.0/93.2/96.8 | 64.8//89.1 | /94.7/96.9 | /81.8/94.3 | 79.4/96.0/98.0 |
| Zipper | 79.7/96.0/99.1 | 62.8/95.8/98.9 | 81.5/96.8/98.8 | 57.5/95.3/98.3 | 76.4/95.2/98.4 | 56.0/95.7/98.6 | 50.1/93.8/97.9 | 60.0/91.2/97.8 | 85.2//99.1 | /95.4/96.6 | /96.3/98.8 | 90.2/98.9/99.7 |
| Average | 75.7/95.0/98.9 | 61.1/94.2/98.4 | 66.1/89.2/97.0 | 60.8/94.4/97.9 | 64.0/89.3/96.0 | 61.5/94.5/98.1 | 63.1/93.4/98.0 | 62.7/92.0/98.2 | 75.6//97.9 | /95.2/96.6 | /94.5/98.4 | 81.3/97.2/99.3 |
| Total Average | 74.2/95.3/99.0 | 57.1/93.1/98.1 | 68.4/91.3/97.3 | 55.4/93.9/97.8 | 65.5/90.8/96.3 | 62.5/95.2/98.2 | 59.8/93.3/97.9 | 58.1/91.1/97.7 | 75.8//97.9 | /95.9/96.8 | /95.0/98.3 | 81.2/97.5/99.3 |
| Category | PRN | BGAD | DevNet | DRA | Ours (Sup) | Ours (Semi) |
| (Zhang et al., 2023a) | (Yao et al., 2023) | (Pang et al., 2021) | (Ding et al., 2022) | |||
| Carpet | 82.0/97.0/99.0 | 83.2/98.9/99.6 | 45.7/85.8/97.2 | 52.3/92.2/98.2 | 89.1/99.1/99.7 | 88.9/99.1/99.7 |
| Grid | 45.7/95.9/98.4 | 59.2/98.7/98.4 | 25.5/79.8/87.9 | 26.8/71.5/86.0 | 66.4/97.0/99.4 | 71.5/98.5/99.7 |
| Leather | 69.7/99.2/99.7 | 75.5/99.5/99.8 | 8.1/88.5/94.2 | 5.6/84.0/93.8 | 81.7/99.7/99.9 | 82.0/99.6/99.9 |
| Tile | 96.5/98.2/99.6 | 94.0/97.9/99.3 | 52.3/78.9/92.7 | 57.6/81.5/92.3 | 96.9/98.9/99.7 | 96.6/98.7/99.7 |
| Wood | 82.6/95.9/97.8 | 78.7/96.8/98.0 | 25.1/75.4/86.4 | 22.7/69.7/82.9 | 88.7/97.9/99.2 | 86.2/97.1/98.6 |
| Average | 75.3/97.2/98.9 | 78.1/98.4/99.2 | 31.3/81.7/91.7 | 33.0/79.8/90.6 | 84.7/98.5/99.5 | 85.0/98.6/99.5 |
| Bottle | 92.3/97.0/99.4 | 87.1/97.1/99.3 | 51.5/83.5/93.9 | 41.2/77.6/91.3 | 93.6/98.5/99.5 | 93.6/98.4/99.5 |
| Cable | 78.9/97.2/98.8 | 81.4/97.7/98.5 | 36.0/80.9/88.8 | 34.7/77.7/86.6 | 89.5/95.9/99.2 | 86.5/96.3/99.3 |
| Capsule | 62.2/92.5/98.5 | 58.3/96.8/98.8 | 15.5/83.6/91.8 | 11.7/79.1/89.3 | 60.0/97.0/98.8 | 58.4/97.6/99.1 |
| Hazelnut | 93.8/97.4/99.7 | 82.4/98.6/99.4 | 22.1/83.6/91.1 | 22.5/86.9/89.6 | 92.2/98.3/99.8 | 86.0/97.3/99.7 |
| Metal nut | 98.0/95.8/99.7 | 97.3/96.8/99.6 | 35.6/76.9/77.8 | 29.9/76.7/79.5 | 99.1/98.2/99.9 | 98.3/98.1/99.8 |
| Pill | 91.3/97.2/99.5 | 92.1/98.7/99.5 | 14.6/69.2/82.6 | 21.6/77.0/84.5 | 86.1/98.9/99.3 | 89.6/98.9/99.5 |
| Screw | 44.9/92.4/97.5 | 55.3/96.8/99.3 | 1.4/31.1/60.3 | 5.0/30.1/54.0 | 72.1/98.8/99.8 | 67.9/98.6/99.8 |
| Toothbrush | 78.1/95.6/99.6 | 71.3/96.4/99.5 | 6.7/33.5/84.6 | 4.5/56.1/75.5 | 74.2/97.1/99.6 | 73.3/96.7/99.6 |
| Transistor | 85.6/94.8/98.4 | 82.3/97.1/97.9 | 6.4/39.1/56.0 | 11.0/49.0/79.1 | 85.5/97.8/98.6 | 86.4/97.9/98.6 |
| Zipper | 77.6/95.5/98.8 | 78.2/97.7/99.3 | 19.6/81.3/93.7 | 42.9/91.0/96.9 | 91.0/99.2/99.7 | 91.3/99.2/99.8 |
| Average | 80.3/95.5/99.0 | 78.6/97.4/99.1 | 20.9/66.3/82.1 | 22.5/70.1/82.6 | 84.3/98.0/99.4 | 83.2/97.9/99.5 |
| Total Average | 78.6/96.1/99.0 | 78.4/97.7/99.2 | 24.4/71.4/85.3 | 26.0/73.3/85.3 | 84.4/98.1/99.5 | 83.8/98.1/99.5 |
3.5 Results on BTAD
As a more challenging alternative to MVTec-AD, BTAD (Mishra et al., 2021) (beanTech Anomaly Detection) contains high-resolution color images of three industrial products. Each product includes normal images in the train set and the corresponding test set consists of both defective and defect-free images.
We further evaluate our algorithm on the BTAD dataset with those SOTA methods also reporting their results on this dataset. Table 3 shows that SemiREST achieves comparable performances to the unsupervised SOTA. Furthermore, as shown in Table 4, with full supervision, the proposed method surpasses SOTA methods by a large margin (, and ) for all three metrics. Similarly to the situation of MVTec-AD, the semi-supervised SemiREST also obtains higher average performances than the supervised SOTA algorithms.
| Category | PatchCore | DRAEM | SSPCAB | CFLOW | RD | PyramidFlow | NFAD | RD++ | Ours (Un) |
| (Roth et al., 2022) | (Zavrtanik et al., 2021) | (Ristea et al., 2022) | (Gudovskiy et al., 2022) | (Deng and Li, 2022) | (Lei et al., 2023) | (Yao et al., 2023) | (Tien et al., 2023) | ||
| 01 | 47.1/78.4/96.5 | 17.0/61.4/91.5 | 18.1/62.8/92.4 | 39.6/60.1/94.8 | 49.3/72.8/95.7 | //97.4 | 46.7/76.6/96.7 | /73.2/96.2 | 52.4/83.9/97.5 |
| 02 | 56.3/54.0/94.9 | 23.3/39.0/73.4 | 15.8/28.6/65.6 | 65.5/56.9/93.9 | 66.1/55.8/96.0 | //97.6 | 59.2/57.9/96.4 | /71.3/96.4 | 63.1/61.5/96.5 |
| 03 | 51.2/96.4/99.2 | 17.2/84.3/96.3 | 5.0/71.0/92.4 | 56.8/97.9/99.5 | 45.1/98.8/99.0 | //98.1 | 62.8/98.8/99.7 | /87.4/99.7 | 50.9/98.8/99.7 |
| Average | 51.5/76.3/96.9 | 19.2/61.6/87.1 | 13.0/54.1/83.5 | 54.0/71.6/96.1 | 53.5/75.8/96.9 | //97.7 | 56.2/77.8/97.6 | /77.3/97.4 | 55.5/81.4/97.9 |
| Category | BGAD | PRN | Ours (Sup) | Ours (Semi) |
| (Yao et al., 2023) | (Zhang et al., 2023a) | |||
| 01 | 64.0/86.6/98.4 | 38.8/81.4/96.6 | 81.3/93.1/99.0 | 69.8/86.2/98.2 |
| 02 | 83.4/66.5/97.9 | 65.7/54.4/95.1 | 84.7/81.4/98.1 | 81.2/69.0/97.9 |
| 03 | 77.4/99.5/99.9 | 57.4/98.3/99.6 | 79.9/99.4/99.9 | 75.6/99.6/99.9 |
| Average | 74.9/84.2/98.7 | 54.0/78.0/97.1 | 82.0/91.3/99.0 | 75.5/84.9/98.7 |
3.6 Results on KolektorSDD2
KolektorSDD2 (Božič et al., 2021) dataset is designed for surface defect detection and includes various types of defects, such as scratches, minor spots, and surface imperfections. It comprises a training set with positive (defective) and negative (defect-free) images, as well as a test set with positive and negative images. We compare the performances of SemiREST with the SOTA results available in the literature. In the unsupervised setting, the algorithms are tested on the original test set. For the supervised and semi-supervised settings, a new training set is generated by combining all the normal training images and defective images which are randomly selected from the original training set.
As shown in Table 5, Our unsupervised performances beat those of SOTA methods by a large margin (, and for AP, PRO and Pixel-AUROC, respectively). Under supervised and semi-supervised settings, our method also achieves better results. It is worth noting that the unsupervised SemiREST performs better than itself with full supervision and semi-supervision. This over-fitting phenomenon might be caused by the (unnecessarily) low sampling rate of defective images in training.
| Method | AP | PRO | AUROC |
| PatchCore (Roth et al., 2022) | 64.1 | 88.8 | 97.1 |
| DRAEM (Zavrtanik et al., 2021) | 39.1 | 67.9 | 85.6 |
| SSPCAB (Ristea et al., 2022) | 44.5 | 66.1 | 86.2 |
| CFLOW (Gudovskiy et al., 2022) | 46.0 | 93.8 | 97.4 |
| RD (Deng and Li, 2022) | 43.5 | 94.7 | 97.6 |
| Ours (Un) | 72.8 | 98.1 | 99.4 |
| PRN (Zhang et al., 2023a) | 72.5 | 94.9 | 97.6 |
| Ours (Sup) | 73.6 | 96.7 | 98.0 |
| Ours (Semi) | 72.1 | 97.5 | 99.1 |
3.7 Analysis on weak labels
Recall that the main motivation of this paper is to reduce the labeling cost of AD tasks, we report the annotation time-consumption of the proposed two weak labels compared with pixel-level annotations.
To obtain the labeling time, the pixel labels, block labels and the bounding boxes of anomaly regions on a subset of MVTec-AD ( defective images for each sub-category) are all manually annotated. Four master students majoring in computer vision complete the labeling task using a self-developed labeling tool, as shown in Figure 2. The annotators are asked to mimic the ground truth annotations shown in a sub-window of the GUI and the annotating time for each image is recorded.
The average annotation times of three kinds of labels are illustrated in Figure 3, along with the corresponding best AD performances (Pixel-AUROC, PRO, AP). According to the figure, one requires only around seconds for labeling bounding boxes and around seconds for generating block-wise labels, on one image. In contrast, the pixel-label consumes more than seconds for one image, while with consistently lower accuracy.
3.8 Comparison of model size and speed
We perform a comparison of model size and speed with some typical SOTA methods and report the results in Table 6. From the table, we can see that the sliding window strategy drags down the speed of SemiREST to some extent. However, the speed increases by times when we remove this process with a single Swin transformer model, at the cost of the AP drop.
| Method | AP (%) | Parameters (M) | Latency (ms) | FPS |
| DeSTSeg | 75.8 | 35.15 | 9.3 | 107.5 |
| NFAD | 62.5 | 41.78 | 10.4 | 96.2 |
| DRAEM | 68.4 | 69.05+28.37 | 13.0+6.4 | 51.5 |
| DMAD | 59.8 | 68.57+24.92 | 9.8+20.1 | 34.8 |
| Ours⋆ | 80.3 | 49.73+51.44 | 28.3+15.9 | 22.6 |
| Ours | 81.2 | 49.73+51.44 | 28.3+91.4 | 8.4 |
3.9 Ablation study
In this section, the most influential modules of SemiREST are evaluated in the manner of an ablation study. The involved modules include: the usage of the PCF Section LABEL:subsubsec:pcf, the Bagging prediction of Swin Transformers Section LABEL:subsubsec:bagging, the -NN augmentation of PCRs Section LABEL:subsubsec:knn, the customized MixMatch algorithm Section LABEL:subsec:mixmatch and the random dropout scheme Section LABEL:subsubsec:dropout, respectively. From Table 7 one can observe a consistent increase in the AD performance as more modules are added to the SemiREST model.
In addition, two element-wise distance functions (see Section LABEL:subsubsec:pcf), i.e., the absolute value of the difference (ABS) and the difference square (Square) are also compared in Table 7. According to the table, the square function outperforms the ABS function within unsupervised and semi-supervised settings where no real defective pixels are seen or labeled. The performance gain might be related to the “feature-selection” property of the square operation, which only focuses on the significant components of the residual vector. On the contrary, when the real defective samples are given for generating PCRs, more useful information can be maintained via the conservative ABS function.
Finally, the impact of the backbone model (Swin Transformer, ViT, and the Unet model of DRAEM (Zavrtanik et al., 2021)) and the block size are also analyzed in Table 7. One can see that the combination of Swin transformer and blocks achieves the best performance. Conventional ViT slightly drags down the AD accuracy while the Unet leads to much poorer performances.
| Module | Performance | |||||||||
| Backbone | Block | PCF | Bagging | -NN Augmentation | MixMatch | Random Dropout | Unsupervised | Supervised | Semi-supervised | |
| Square | Swin | 70.5/95.4/98.4 | 75.1/96.6/98.9 | – | ||||||
| ABS | Swin | 69.3/94.8/98.3 | 76.8/96.9/99.0 | – | ||||||
| Square | Swin | 72.4/95.8/98.9 | 76.4/96.6/99.0 | – | ||||||
| Square | Swin | 79.3/96.9/98.8 | 83.4/98.2/99.4 | – | ||||||
| Square | Swin | 81.2/97.5/99.3 | 84.4/98.2/99.5 | 80.7/97.5/99.3 | ||||||
| Square | Swin | 81.2/97.5/99.3 | 84.4/98.2/99.5 | 82.1/97.8/99.4 | ||||||
| Square | Swin | 81.2/97.5/99.3 | 84.4/98.2/99.5 | 83.8/98.1/99.5 | ||||||
| Square | ViT | 80.5/97.1/99.2 | 82.4/97.9/99.1 | 82.9/97.9/99.4 | ||||||
| Square | Unet | 76.9/95.5/98.6 | 84.3/97.8/99.3 | 76.4/94.0/97.1 | ||||||
| Square | Swin | 80.2/97.2/99.2 | 83.6/98.0/99.4 | 83.0/97.9/99.5 | ||||||
| Square | Swin | 75.5/95.8/99.0 | 77.5/96.7/99.3 | 76.9/96.5/99.3 | ||||||
| ABS | Swin | 76.4/96.3/98.7 | 84.9/98.3/99.5 | 83.2/98.0/99.3 | ||||||
3.10 Data availability statement
The datasets used in this study are publicly available from the following sources:
MVTec-AD (Bergmann et al., 2019) dataset can be downloaded from https://www.mvtec.com/company/research/datasets/mvtec-ad.
BTAD(Mishra et al., 2021) dataset can be found at http://avires.dimi.uniud.it/papers/btad/btad.zip.
KolektorSDD2 (Božič et al., 2021) dataset can be found at https://www.vicos.si/resources/
kolektorsdd2/.
4 Conclusion
In this paper, we propose to solve the AD problem via block-wise classifications which require much less annotation effort than pixel-wise segmentation. To achieve this, a sliding vision transformer is employed to predict block labels, based on the smartly designed position-constrained residuals. The proposed bagging strategy of the Swin Transformers lead to the new SOTA accuracy on three well-known AD datasets. In addition, even cheaper bounding-box labels are proposed to further reduce the labeling time. Given only partially labeled normal regions, the ResMixMatch learning scheme successfully exploits the information of unlabeled regions and achieves the AD performances close to that with full-supervisions. The proposed SemiREST algorithm brings record-breaking AD performances to the literature while only requiring much coarser annotations or in short, our SemiREST is cheaper in annotation and better in accuracy.
Thus, SemiREST paves a novel way to reduce the annotation cost for AD problems while maintaining accuracy. According to the experiment of this work, the weak/semi-supervised setting seems a more practical alternative to the classic few-shot setting that directly limits the number of training images. In the future, we believe that better semi-supervised AD algorithms will be developed by exploiting more useful information from the unlabeled image regions.
Acknowledgements: This work was supported by National Key R&D Program of China (No. 2022ZD0118700) and the Fund for Less Developed Regions of the National Natural Science Foundation of China (Grant No. 61962027).